# Dual-Personalizing Adapter for Federated Foundation Models

### v1

@@ Abstract
Recently, foundation models, particularly large language models (LLMs), have demonstrated an impressive ability to adapt to various tasks by fine-tuning large amounts of instruction data. Notably, federated foundation models emerge as a privacy preservation method to fine-tune models collaboratively under federated learning (FL) settings by leveraging many distributed datasets with non-IID data. To alleviate communication and computation overhead, parameter-efficient methods are introduced for efficiency, and some research adapted personalization methods to federated foundation models for better user preferences alignment. However, a critical gap in existing research is the neglect of test-time distribution shifts in real-world applications. Therefore, to bridge this gap, we propose a new setting, termed test-time personalization, which not only concentrates on the targeted local task but also extends to other tasks that exhibit test-time distribution shifts. To address challenges in this new setting, we explore a simple yet effective solution to learn a comprehensive foundation model. Specifically, a dual-personalizing adapter architecture (FedDPA) is proposed, comprising a global adapter and a local adapter for addressing test-time distribution shifts and personalization, respectively. Additionally, we introduce an instance-wise dynamic weighting mechanism to optimize the balance between the global and local adapters, enhancing overall performance. The effectiveness of the proposed method has been evaluated on benchmark datasets across different NLP tasks.
@@ 1 Introduction
Recently, foundational models, especially the large language model (LLM) within the domain of natural language processing (NLP), have garnered significant interest Brown et al. (2020). By utilizing vast amounts of data and sophisticated training algorithms, foundation models are endowed with a rich tapestry of generalized knowledge. To further refine these models for a more precise alignment with specific tasks and user preferences, various fine-tuning methods have been explored. Notably, federated foundation models Zhuang et al. (2023); Yu et al. (2023) represent an innovative approach, integrating federated learning frameworks for collaboratively fine-tuning the pre-trained foundation models by leveraging client-specific datasets to address privacy concerns.
Notwithstanding their considerable potential, direct fine-tuning of foundation models in FL incurs substantial computational and communication overhead due to the voluminous number of parameters encompassed. Parameter-efficient fine-tuning (PEFT) methods Xu et al. (2023a) present a promising alternative to mitigate these challenges. By selectively tuning and transmitting only a subset of parameters in FL, these methods seek to enhance efficiency. Among these PEFT methods, the adapter family Hu et al. (2023) (e.g., LoRA Hu et al. (2021)) stands out as one of the most popular methods in contemporary research of federated foundation models due to its flexibility compared with other methods.
Existing works have explored a spectrum of methodologies to address specific challenges when adapting the adapter-based PEFT methods within federated foundation models, including data heterogeneity Babakniya et al. (2023); Jiang et al. (2023), communication overheads Xu et al. (2023b); Sun et al. (2023), and so on. However, there are few studies delving into the personalization of federated foundation models for local clients, a crucial aspect in practical scenarios characterized by non-uniform data distribution across clients and the presence of diverse ability preferences among different clients. Furthermore, conventional personalized FL approaches Tan et al. (2022) predominantly concentrate on the specific targeted task for each client, often neglecting the challenges posed by test-time distribution shift when sometimes clients encounter other tasks during testing. For example, a client focusing on paper writing may sometimes need translation as an assistant. Given that a global model, trained on all mixed data, could not consistently excel in specifically targeted abilities as highlighted in Wang et al. (2023), there is a pressing need to devise methods for training personalized models that not only prioritize the performance on targeted abilities but also ensure satisfactory outcomes across other tasks, namely distribution shift on deploying federated foundation models.
To solve the aforementioned challenge, we introduce a brand-new setting close to real-world applications, noted as test-time personalization, which follows: 1) each client needs to train a personalized model using its own data from a target task, and 2) during testing, each client’s personalized model needs to be capable to tackle the receiving new tasks (unseen in training) with different distributions (test-time distribution shift). This setting considers the test-time distribution shift scenario rather than only targeted tasks for personalization, which is essential in real applications of foundation models when clients may require assistance with tasks beyond their primary target task. Under this setting, a test-time personalized model should perform well on the personalized targeted task and perform comparably to other test-time tasks. To make this setting more challenging, we consider an extreme scenario where the proprietary tasks barely overlap across clients in training, but during the test, we consider test-time personalization in an ideal setting, where all tasks are included in all clients. As such, we take tasks of other clients with different distributions as test-time tasks for each client.
@@
In test-time personalization, challenges become how to learn general knowledge from different clients for test-time tasks while prioritizing targeted ability for personalization. To address this issue, we explore a simple yet effective method, dubbed dual-personalizing adapter (FedDPA), where each client learns a global adapter to learn general knowledge from the aggregation for test-time tasks and maintains a local adapter for targeted ability personalization. During the inference phase, the local and global adapters are synergistically integrated to facilitate prediction, and an instance-wise dynamic weighting mechanism is proposed to autonomously adjudicate the proportional contribution of the local and global adapters. Experimental results demonstrate that our method achieves state-of-the-art performance on benchmarks. Our main contributions are summarized as follows:
  • We propose a novel setting in personalized federated learning that emphasizes test-time distribution shifts in practical application scenarios, promoting comprehensive model learning during testing.
  • We present a new method to realize test-time personalization, emphasizing the learning of both general and personalized knowledge for a more comprehensive model on various tasks.
@@
  • We conduct an exhaustive analysis using heterogeneous FL benchmarks across diverse NLP tasks. The empirical outcomes reveal that our method attains state-of-the-art performance, underscoring its superior test-time personalization capabilities in contrast to existing methods.
Figure 1: The overall framework of our proposed method. Each client contains a frozen LLM, a trainable global adapter (LoRA) and a trainable local adapter (LoRA) with a specific task, where the global adapter (LoRA) is for test-time tasks and the local adapter (LoRA) is for personalization. During the training, only the parameters of the global adapter (LoRA) are transmitted to the server for aggregation.
@@ 2.1 Adapter-based PEFT Methods
Given the substantial computational and storage burdens associated with directly fine-tuning foundation models, the community has shifted towards embracing parameter-efficient methods Xu et al. (2023a), with the adapter family Hu et al. (2023) being a notable exemplar. According to different architectures, methods in the adapter family can be categorized into four types. The first one is prompt-based learning Lester et al. (2021); Li and Liang (2021), which is aimed at learning the continuous/soft prompt for discrete optimization. The second one is reparametrization-based methods Hu et al. (2021); Edalati et al. (2022), achieving parameter efficiency by utilizing low-rank techniques to decompose the high-dimensional matrices. The third one is series Adapters Houlsby et al. (2019), which introduce additional learnable modules in a sequential manner within specific sublayers. The last one is parallel Adapters He et al. (2021), which focus on learning additional learnable modules in a parallel way with distinct sublayers. In this context, our exploration delves into the adapter-based PEFT methods of federated foundation models.
@@ 2.2 Federated Foundation Models
With the advent of foundation models, there has been a burgeoning interest Zhuang et al. (2023); Yu et al. (2023) in integrating these models within the FL setting. Particularly, in light of the inherent computation and communication cost, recent work Kuang et al. (2023); Zhang et al. (2023b) endeavors have delved deeper into integrating adapter-based parameter-efficient tuning (PEFT) methods with federated foundation models. Building upon this, a multitude of studies have emerged to navigate the challenges of incorporating federated foundation models with adapter-based PEFT methods. The paper Zhang et al. (2023a) stands at the forefront, initiating the integration of instruction tuning within federated LLM frameworks. Addressing data-related issues, the paper Babakniya et al. (2023) introduced a data-driven initialization approach to mitigate the primary challenges associated with LoRA in highly heterogeneous data scenarios. In addition, the research presented in Jiang et al. (2023) proposed a method to annotate unlabeled client-side data by harnessing the prowess of large models to address data scarcity concerns. To further optimize the communication and computational overheads associated with federated foundation models, the works Xu et al. (2023b); Sun et al. (2023) emphasize advancing gradient-free optimization methods suitable for devices with limited memory and computing power. For personalization, paper Yi et al. (2023) focused on designing a specific training paradigm for LoRA to achieve more effective personalization in visual model-heterogeneous scenarios. Diverging from these approaches, our work delves into the realm of personalization with adapters in federated foundation models, extending the scope of research in this area.
@@ 2.3 Personalized Federated Learning
To address the necessity of personalization for individual clients, personalized Federated Learning (PFL) Tan et al. (2022), which aims at training to cater to individual client preferences and needs, is proposed. Broadly, existing PFL methods can be categorized into two primary types: fine-tuning the global model for personalization or learning additional personalized models. Research works Fallah et al. (2020); Collins et al. (2021) in the first category fine-tuned the whole or part of the global model with each client’s local dataset for personalization. While research works Li et al. (2021a, b) in the second category is to learn the additional personalized layers or model through local aggregation. Nonetheless, a prevalent limitation among these PFL approaches is their concentrated focus on a specifically targeted task, often at the expense of performance when encountering test-time distribution shifts. Such a constrained approach might be suboptimal for practical applications that demand capability across other tasks. Therefore, we introduce the concept of test-time personalization to fill this gap.
@@ 3 Problem Definition
3.1 General FL Framework
In this paper, our method is built upon the general FL framework. In this framework, there are $M$ clients, each possessing its distinct local privacy data ${\mathbb{D}}^{m}$, where $m$ indexes a client. Concurrently, a server at the center level oversees the learning and maintenance of the aggregated model. For model aggregation across clients, we employ the FedAvg algorithm McMahan et al. (2017) without sacrificing generality, where the aggregated parameters ${\bm{\theta}}$ are derived from a weighted sum of all clients’ parameters ${\bm{\theta}}^{m}$ based on the numbers of data. However, to address potential biases stemming from different numbers of tasks and data privacy concerns, our model aggregation strategy is based on the client number rather than the number of data. Consequently, the formulation can be represented as:
${\bm{\theta}}=\sum^{M}_{i=1}\frac{1}{M}{\bm{\theta}}^{i}$ (1)
Local Tasks.
In general FL framework, each local dataset, denoted as ${\mathbb{D}}^{m}$, is characterized by a set of data pairs represented as ${\{(x_{i}^{m},y_{i}^{m})|i\in{1,...,N^{m}}\}}$, with $N^{m}$ denoting the number of data. The objective for each local client is to minimize the empirical risk of all local data, which can be formulated as
$\min_{{\bm{\theta}}^{m}}\sum_{(x,y)\in{\mathbb{D}}^{m}}\mathcal{L}(x,y;{\bm{\theta}}^{m})$ (2)
where $\mathcal{L}$ is the loss function applied to each data instance and parameterized by ${\bm{\theta}}^{m}$.
3.2 Test-time Personalization
Here, we will introduce the training and testing phases separately for our test-time personalization setting.
@@ 3.2 Test-time Personalization
In the training phase, the process and objective are the same as the general FL framework, where the model trains on the local data for empirical risk minimization. While in the testing phase, except for the test set ${\mathbb{D}}_{s}$ driven from the same domain as training data, there are also some test sets ${\mathbb{D}}_{t}$ under data distribution shifts $p_{s}(x,y)\neq p_{t}(x,y)$, and we call these datasets ${\mathbb{D}}_{t}$ as test-time datasets. Therefore, the objective of the model should not only perform well on the test set ${\mathbb{D}}_{s}$ but also have comparable results on the test-time dataset ${\mathbb{D}}_{t}$. This objective is consistent with the practical scenarios, since users primarily focus on the abilities they often utilize (abundant data available for training) and occasionally also introduce new tasks (limited to test data).

### v2

@@ Abstract
Recently, foundation models, particularly large language models (LLMs), have demonstrated an impressive ability to adapt to various tasks by fine-tuning diverse instruction data. Notably, federated foundation models (FedFM) emerge as a privacy preservation method to fine-tune models collaboratively under federated learning (FL) settings by leveraging many distributed datasets with non-IID data. To alleviate communication and computation overhead, parameter-efficient methods are introduced for efficiency, and some research adapted personalization methods to FedFM for better user preferences alignment. However, a critical gap in existing research is the neglect of test-time distribution shifts in real-world applications, and conventional methods for test-time distribution shifts in personalized FL are less effective for FedFM due to their failure to adapt to complex distribution shift scenarios and the requirement to train all parameters. To bridge this gap, we refine the setting in FedFM, termed test-time personalization, which aims to learn personalized federated foundation models on clients while effectively handling test-time distribution shifts simultaneously. To address challenges in this setting, we explore a simple yet effective solution, a Federated Dual-Personalizing Adapter (FedDPA) architecture. By co-working with a foundation model, a global adapter and a local adapter jointly tackle the test-time distribution shifts and client-specific personalization. Additionally, we introduce an instance-wise dynamic weighting mechanism that dynamically integrates the global and local adapters for each test instance during inference, facilitating effective test-time personalization. The effectiveness of the proposed method has been evaluated on benchmark datasets across different NLP tasks with released code.
@@ 1 Introduction
Foundation models, especially the large language model (LLM) in natural language processing (NLP), have nearly exhausted public data sources for training. This necessitates alternative solutions to further improve these foundation models by leveraging private or protected data sources, such as business data in companies, smartphones, and so on. Federated foundation models (FedFM)zhuang2023foundation ; yu2023federated offer a promising solution by integrating federated learning (FL) frameworks to enhance the foundation models in a decentralized manner. Built upon existing Parameter-efficient fine-tuning (PEFT) methods xu2023parameter ; hu2023llm ; hu2021lora , FedFM is a collaboratively fine-tuning framework that leverages private datasets with privacy preservation and avoiding overfitting from client-specific fine-tuning.
Test-time distribution shift in FL jiang2023test ; tan2024heterogeneity presents a significant challenge in practical scenarios, as clients may encounter unseen learning tasks during the testing and model inference phases. For example, a client accustomed to writing emails in English may require translation assistance when working on a new project in Chinese. Therefore, it is imperative for the deployed machine learning model to be capable of tackling the test-time distribution shifts from the training data, and our paper addresses this critical issue of test-time distribution shifts within the FedFM scenario. Previous works in test-time FLtan2024heterogeneity ; jiang2023test predominantly utilize conventional deep learning models that are trained from scratch in federated settings. Recent FedFM methods mainly focus on addressing specific challenges related to data heterogeneity babakniya2023slora ; jiang2023low and communication overheads xu2023federated ; sun2023fedbpt . However, none of these methods have discussed test-time distribution shifts in FedFM scenarios.
To fill this gap, we propose a novel FedFM framework that is robust to client-specific alignment and test-time distribution shifts simultaneously. With the support of a foundation model with PEFT methods, we first refine the federated setting, termed test-time personalization, which follows: 1) each client needs to train a personalized model using its own data from a target task, and 2) during testing, each client’s personalized model needs to be robust to tackle the receiving new tasks (unseen in training) with different distributions (test-time distribution shift). Essentially, the proposed test-time personalization in FL could be simply viewed as an optimization task to seek a sweet trade-off between client-specific model personalization and model generalization to test data.
@@
For test-time personalization in FedFM, two primary challenges—test-time distribution shifts and personalization—necessitate learning tailored to distinct objectives, and the training cost of foundation models also represents a significant concern. To address these issues, we explore a simple yet effective method, dubbed Federated Dual-Personalizing Adapter (FedDPA), where each client learns a global adapter to learn generic knowledge from the aggregation for test-time tasks and maintains a local adapter for targeted ability personalization. During the inference phase, the local and global adapters are dynamically integrated to facilitate prediction, where an instance-wise dynamic weighting mechanism is proposed to autonomously adjudicate the proportional contribution of the local and global adapters for each test instance. Experimental results demonstrate that our method achieves state-of-the-art performance on benchmarks and all data and code are released  [^1] https://github.com/Lydia-yang/FedDPA. Our main contributions are summarized as follows:
  • We are the first to explore the test-time distribution shifts problem in federated foundation models for practical application scenarios alignment.
  • We introduce a new method, namely dual-personalizing adapter, to emphasize learning both generic and personalized knowledge in the context of FedFM with test-time personalization.
@@
  • We conduct an exhaustive analysis using heterogeneous FL benchmarks across diverse NLP tasks. The empirical outcomes reveal that our method attains state-of-the-art performance, underscoring its superior test-time personalization capabilities than existing methods.
@@ 2.1 Adapter-based PEFT Methods
Given the substantial computational and storage burdens associated with directly fine-tuning foundation models, the community has shifted towards embracing parameter-efficient methods xu2023parameter , with the adapter family hu2023llm being a notable exemplar. According to different architectures, methods in the adapter family can be categorized into four types. The first one is prompt-based learning lester2021power ; li2021prefix , which is aimed at learning the continuous/soft prompt for discrete optimization. The second one is reparametrization-based methods hu2021lora ; edalati2022krona , achieving parameter efficiency by utilizing low-rank techniques to decompose the high-dimensional matrices. The third one is series Adapters houlsby2019parameter , which introduce additional learnable modules in a sequential manner within specific sublayers. The last one is parallel Adapters he2021towards , which focus on learning additional learnable modules in a parallel way with distinct sublayers. In this context, our exploration delves into the adapter-based PEFT methods of federated foundation models.
@@ 2.2 Federated Foundation Models
With the advent of foundation models, there has been a burgeoning interest zhuang2023foundation ; yu2023federated ; ren2024advances ; charles2024towards in integrating these models within the FL setting. Particularly, in light of the inherent computation and communication cost, recent work kuang2023federatedscope ; zhang2023fedpetuning ; chen2024feddat endeavors have delved deeper into integrating adapter-based parameter-efficient tuning (PEFT) methods with federated foundation models. Building upon this, a multitude of studies have emerged to navigate the challenges of incorporating federated foundation models with adapter-based PEFT methods. The paper zhang2023towards stands at the forefront, initiating the integration of instruction tuning within federated LLM frameworks. Addressing data-related issues, the paper babakniya2023slora introduced a data-driven initialization approach to mitigate the primary challenges associated with LoRA in highly heterogeneous data scenarios. In addition, the research presented in jiang2023low proposed a method to annotate unlabeled client-side data by harnessing the prowess of large models to address data scarcity concerns. To further optimize the communication and computational overheads associated with federated foundation models, the works xu2023federated ; sun2023fedbpt ; xu2024fwdllm emphasize advancing gradient-free optimization methods suitable for devices with limited memory and computing power. For personalization, paper yi2023fedlora focused on designing a specific training paradigm for LoRA to achieve more effective personalization in visual model-heterogeneous scenarios. Diverging from these approaches, our work delves into the realm of personalization with adapters in federated foundation models, extending the scope of research in this area.
@@ 2.3 Personalized Federated Learning
To address the necessity of personalization for individual clients, personalized Federated Learning (PFL) tan2022towards , which aims at training to cater to individual client preferences and needs, is proposed. Broadly, existing PFL methods can be categorized into two primary types: fine-tuning the global model for personalization or learning additional personalized models. Research works fallah2020personalized ; collins2021exploiting in the first category fine-tuned the whole or part of the global model with each client’s local dataset for personalization. While research works li2021ditto ; li2021fedbn in the second category is to learn the additional personalized layers or model through local aggregation. Nonetheless, a prevalent limitation among these PFL approaches is their concentrated focus on a specifically targeted task, often at the expense of performance when encountering test-time distribution shifts.
To fill this gap, recent research has shifted focus towards exploring different test-time distribution shifts in PFL. In contrast to studies yoon2021federated ; dong2022federated in federated continual/incremental learning possessing ample annotated data from different distributions for training to address shifts, test-time FL focuses on handling distribution shifts during testing without the availability of annotated data for further training. One strand of research bao2024adaptive ; xu2023joint concentrates on addressing test-time distribution shifts that occur when new clients are introduced during the testing phase by module/prior adaptation. Another line of studiestan2024heterogeneity ; jiang2023test aims to tackle distribution shifts in existing clients during testing by aligning test features with existing features. Our paper falls into the second type and differs from previous work by exploring this challenge within the framework of foundation models, which are characterized by extensive parameter scales and more complex test-time distribution shifts.
Figure 1: The overall framework of FedDPA. Each client contains a frozen LLM, a trainable global adapter (LoRA) and a trainable local adapter (LoRA) with a specific task, where the global adapter (LoRA) is for test-time tasks and the local adapter (LoRA) is for personalization. During the training, only the parameters of the global adapter (LoRA) are transmitted to the server for aggregation.
@@ 3 Problem Definition
3.1 Test-time Personalization in FedFM
Considering $M$ clients in an FL system, each client possesses its distinct local training dataset ${\mathbb{D}}^{m}_{train}$ and test dataset ${\mathbb{D}}^{m}_{test}$, where $m$ indexes a client. One data pair in datasets is denoted as ${({\bm{x}},{\bm{y}})}$, where ${\bm{x}}$ is the input data and ${\bm{y}}$ is its corresponding label.
In each client $m$, we will introduce the training and testing phases separately for our test-time personalization setting. In the training phase, model training utilizes solely the local dataset ${\mathbb{D}}_{train}^{m}$, which is derived from the distribution $P_{s}^{m}$. While in the testing phase, the test dataset ${\mathbb{D}}_{test}^{m}$ comprises two components: the test set ${\mathbb{D}}_{s}^{m}$ driven from the same distribution $P_{s}^{m}$ as training data, and additional test sets ${\mathbb{D}}_{t}^{m}$ under data distribution shifts $P_{s}^{m}({\bm{x}},{\bm{y}})\neq P_{t}^{m}({\bm{x}},{\bm{y}})$. Therefore, the test dataset is ${\mathbb{D}}_{test}^{m}={\mathbb{D}}_{s}^{m}\cup{\mathbb{D}}_{t}^{m}$, and we call these datasets ${\mathbb{D}}_{t}^{m}$ as test-time datasets. Unlike previous workstan2024heterogeneity ; jiang2023test in test-time FL concentrating on either feature-level shifts $P_{s}^{m}({\bm{x}})\neq P_{t}^{m}({\bm{x}})$ or label shifts $P_{s}^{m}({\bm{y}})\neq P_{t}^{m}({\bm{y}})$, we investigate a more complex scenario where various distribution shifts, including semantic shifts, domain shifts and others, exist simultaneously. This is aligned with the practical application of foundation models, which often encounter testing data originating from diverse domains, backgrounds, or populations.
@@ 3.1 Test-time Personalization in FedFM
Therefore, the objective of the model in each client should not only perform well on the test set ${\mathbb{D}}_{s}^{m}$ (refer to personalization) but also have comparable results on the test-time dataset ${\mathbb{D}}_{t}^{m}$ (refer to test-time performance). This objective is consistent with the practical scenarios, since users primarily focus on the abilities they often utilize (abundant data available for training) and occasionally also introduce new tasks (limited to test data).
3.2 Challenges
The test-time personalization setting raises two pivotal considerations: personalization and test-time distribution shifts. Personalization is the primary focus, followed by optimizing test-time tasks. Our proposed method introduced in section 4 is designed to achieve personalization within FedFM while ensuring comparable results for test-time tasks, and its vital intuition is illustrated below.
Considering a foundation model, it comprises a main body $f({\bm{\theta}})$, which holds most of the parameters and processes input ${\bm{x}}$ to produce output features ${\bm{h}}=f({\bm{x}};{\bm{\theta}})$. Additionally, there is a tail $g({\bm{\theta}}_{t})$ that maps these features to the output space (e.g., vocabulary), resulting in the predicted result $\hat{{\bm{y}}}=g({\bm{h}};{\bm{\theta}}_{t})$. Typically, the focus in tuning and adaptations primarily lies on the main body $f({\bm{\theta}})$ because the tail $g({\bm{\theta}}_{t})$, usually a linear function, remains unchanged (frozen) during tuning hu2023llm .
Discordance between Personalization and Test-time Tasks.
The key to addressing test-time distribution shifts lies in learning generic features universally applicable across disparate distributions arjovsky2020out . That is, learning a foundation model $f({\bm{\theta}})$ to satisfy $P_{s}(f({\bm{x}};{\bm{\theta}}),{\bm{y}})=P_{t}(f({\bm{x}};{\bm{\theta}}),{\bm{y}})$ although $P_{s}({\bm{x}},{\bm{y}})\neq P_{t}({\bm{x}},{\bm{y}})$. FL is a methodology designed to learn generic features across diverse non-IID data (different distributions) through aggregation algorithms charles2024towards ; tan2024heterogeneity . Therefore, we tailor FL training for addressing test-time tasks with the objective $\min_{\bm{\theta}}\mathcal{L}_{P_{all}}({{\bm{\theta}}})$, where $\mathcal{L}_{P_{all}}$ represents the loss function designed for learning generic features towards all clients’ distributions $P_{all}$. However, personalization focuses on aligning the model with the specific distribution $P_{s}$, which means learning a foundation model $f({\bm{\theta}})$ with the objective $\min_{\bm{\theta}}\mathcal{L}_{P_{s}}({{\bm{\theta}}})$, where $\mathcal{L}_{P_{s}}$ represents the loss function designed for learning personalized features towards the specific distribution. Therefore, the discordance between specific distribution alignment for personalization and generic feature learning for test-time tasks leads to inconsistent optimization objectives.
The above analysis motivates a dual model strategy—one model for test-time tasks and one model for personalization—to realize test-time personalization in FedFM. This strategy, together with our other techniques presented below for FedFM scenarios, constitutes the foundation of our method.
4 Proposed Method
To align with the application scenarios, we consider the test-time personalization setting in FedFM. Following a similar assumption from a Mixture of Experts masoudnia2014mixture , any test-time task (distributions) to a client can be approximated as a mixture of training tasks seen by other clients in the federated learning system. Therefore, each client primarily personalizes its model based on its local training task, while also tackling unseen test-time tasks by leveraging insights gained from other clients in the federated learning system. Discussions of other scenarios can be found in Appendix C.
In test-time personalization, test-time distribution shifts and personalization are two main issues that need to be addressed, and their optimization objectives toward different distributions are inconsistent. To address these challenges and consider the efficient learning of FedFM, we propose a Federated Dual-Personalizing Adapter (FedDPA) system for each client, as shown in Fig 1. During training, a global adapter is employed to acquire generic features by FL training for test-time tasks (Sec. 4.1). Meanwhile, to address personalization, a local adapter is maintained locally to align with the client’s specific distribution, and leverages generic knowledge from the global adapter for faster learning (Sec. 4.2). During the inference, the learned global and local adapters are dynamically combined using a weight generated by the instance-wise dynamic weighting mechanism for each input test instance, realizing test-time personalization (Sec. 4.3).
The Overall Objective.
Considering the computation and communication cost of FedFM, we utilize the adapter-based PEFT methods, which only learn a small part of parameters $\Delta{\bm{\theta}}$ while keeping most of the parameters ${\bm{\theta}}$ frozen. Our proposed FedDPA is to learn the global adapter $\Delta{\bm{\theta}}_{g}$ and local adapters $\Delta{\bm{\theta}}_{l}^{m}$ simultaneously across $M$ client to realize test-time personalization,
$\min_{\Delta{\bm{\theta}}_{g},\{\Delta{\bm{\theta}}_{l}^{m}\}}$ $\sum_{m=1}^{M}[\mathcal{L}_{({\bm{x}},{\bm{y}})\sim P_{s}^{m}}({\bm{\theta}};\Delta{\bm{\theta}}_{g};\Delta{\bm{\theta}}_{l}^{m})]$ (1)
$s.t.$ $\Delta{\bm{\theta}}_{g}^{*}\in\operatorname*{arg\,min}_{\Delta{\bm{\theta}}_{g}}\mathcal{L}_{({\bm{x}},{\bm{y}})\sim P_{all}}({\bm{\theta}};\Delta{\bm{\theta}}_{g};\Delta{\bm{\theta}}_{l}^{m})$
where the first part is a standard personalized FL loss to find optimal personalized models by minimizing the sum of loss on local training tasks $\mathcal{L}_{({\bm{x}},{\bm{y}})\sim P_{s}^{m}}(.)$, and the second part is a constraint term to seek an optimal solution by minimizing the test-time loss $\mathcal{L}_{({\bm{x}},{\bm{y}})\sim P_{all}(.)}$. Because we assume that the test-time task is unseen to a client but observed by other clients, the test-time loss can be estimated using the Empirical Risk Minimization of all client’s training tasks $\min_{\Delta{\bm{\theta}}_{g}}\sum^{M}_{m=1}r_{m}\mathcal{L}_{m}({\bm{\theta}};\Delta{\bm{\theta}}_{g})$, where $P_{all}$ denotes all distributions of tasks in all clients, $r_{m}$ denotes each client’s weight for aggregation (e.g., in FedAvg, $r_{m}$ is the proportion of each client’s data number to all clients’ data number) and $\mathcal{L}_{m}$ denotes the loss for each client over its local training dataset. Since the above objective cannot be solved directly, we propose to alternatively learn the global and local adapters in a sequential manner (FedDPA-F with local adapter fine-tuning) or iterative manner (FedDPA-T with local adapter training). Detailed algorithms of these two methods are in Appendix A.3.